# Learning Constraints from Stochastic Partially-Observed Closed-Loop Demonstrations

### v1

@@ Abstract
We present an algorithm for learning unknown parametric constraints from locally-optimal input-output trajectory data. We assume that the given data is generated by demonstrators with stochastic nonlinear dynamics who execute a state or output feedback law to robustly satisfy the constraints despite worst-case dynamics and output noise. We encode the Karush-Kuhn-Tucker (KKT) conditions of this robust optimal output feedback control problem within a feasibility problem to recover constraints consistent with the local optimality of the demonstrations. We prove that our constraint learning method (i) accurately recovers the demonstrator’s state or output feedback policy, and (ii) conservatively estimates the set of all state or output feedback policies that ensure constraint satisfaction despite worst-case noise realizations. Moreover, we perform sensitivity analysis, proving that when demonstrations are corrupted by transmission error, the inaccuracy in the learned state or output feedback law scales linearly in the error magnitude. Our method accurately recovers unknown constraints from simulated noisy, closed-loop demonstrations generated using dynamics, both linear and nonlinear, (e.g., unicycle and quadrotor) and a range of state and output feedback mechanisms.
I Introduction
@@
In safety-critical robotics, autonomous agents must infer and enforce hard constraints to guarantee safe operation, such as for collision avoidance in crowded environments. Learning from demonstration (LfD) enables robots to recover unknown parametric constraints from trajectory data [1, 2, 3]. The key idea is that locally-optimal demonstrations generated from a constrained optimal control problem satisfy an associated set of Karush-Kuhn-Tucker (KKT) conditions, which can be inverted to recover constraints via inverse optimal control (IOC). Prior work has shown this is possible for open-loop state and control demonstrations under deterministic dynamics [1, 4]. In practice, however, agents operate with state uncertainty and stochastic dynamics. Thus, instead of open-loop plans, a demonstrator must apply an output feedback control policy to partial, noisy observations. Existing methods cannot recover constraints from such input-output data.
@@
To address this gap, we make the following contributions. We present a novel method to infer a priori unknown parametric constraints from input-output data generated under linear time-varying (LTV) stochastic dynamics with partial observations and output noise. Our method extends the IOC-based constraint learning method in [1] to the partial and noisy observation setting (Sec. III). We evaluate our method on linear and nonlinear (e.g., unicycle, quadrotor) dynamics and a variety of feedback controllers, successfully recovering constraints from the generated demonstrations (Sec. IV).
@@
Our work draws from and contributes to IOC-based LfD for cost and constraint inference. Specifically, [5, 6, 2] introduce IOC methods assuming known constraints, [7, 8] enable cost inference via IOC, while [1, 9, 10] present IOC methods for constraint learning. Unlike [9], which addresses stochastic demonstrations in discrete spaces, our method recovers constraints with provable accuracy guarantees in continuous domains. Compared to [10], which assumes demonstrators with partial obstacle knowledge but full state information, we consider partial state observability. Further, while [1] use deterministic trajectories, we learn constraints from demonstrations generated via output feedback with noisy, partial observations. On a technical level, we leverage system-level synthesis (SLS) [11, 12, 13] to encode output feedback policies for input-output demonstrations. Whereas prior SLS work focuses on controller design, we apply SLS to constraint inference from noisy, partially observed data.
@@
Notation: Given $x_{0},\cdots,x_{T}\in\mathbb{R}^{n}$ and indices $t_{1},t_{2}\in\{0,\ldots,T\}:=[0,T]$ with $t_{1}\leq t_{2}$, let $x_{t_{1}:t_{2}}:=(x_{t_{1}},\cdots,x_{t_{2}})\in\mathbb{R}^{n(t_{2}-t_{1}+1)}$. Given $M_{1},\cdots,M_{K}\in\mathbb{R}^{m\times n}$, we denote by $M:=\text{diag}\{M_{1},\cdots,M_{K}\}\in\mathbb{R}^{mK\times nK}$ the block diagonal matrix with $M_{k}$ as the $k$-th block, $\forall k\in\{1,\ldots,K\}:=[K]$. For any $m,n\in\mathbb{N}$, we denote the $m\times m$ identity matrix by $I_{m}$ and the $m\times n$ zero matrix by $O_{m\times n}$, with subscripts omitted when dimensions are clear from context. Given $A\in\mathbb{R}^{m\times n}$, we denote by $A^{\dagger}\in\mathbb{R}^{n\times m}$ the Moore-Penrose pseudoinverse of $A$.
II Problem Formulation
@@
We formulate the robust optimal control problem used to generate demonstration data (Sec. II-A), present system level synthesis (SLS) for output feedback (Sec. II-B), discuss data generation (Sec. II-C), and state our problem (Sec. II-D).
II-A Robust Optimal Control Problem Formulation
@@
Consider a stochastic dynamical system $f_{t}:\mathbb{R}^{n}\times\mathbb{R}^{n_{i}}\rightarrow\mathbb{R}^{n}$ with additive dynamics and output noise, and initial condition fixed at $x_{0}=\bar{x}$ as shown below: [^1] Uncertainty in $x_{0}$ can be modeled by incorporating it into the definition of the dynamics noise; we do so in our experiments. For details, see [12].
@@
Above, $x_{t}\in\mathbb{R}^{n}$, $u_{t}\in\mathbb{R}^{n_{i}}$, $w_{t}\in\mathbb{R}^{n}$, $y_{t}\in\mathbb{R}^{n_{o}}$, and $e_{t}\in\mathbb{R}^{n_{o}}$ denote the state, input, dynamics noise, output, and output noise at time $t\in[0,T]$, respectively. For convenience, let $\textbf{x}:=x_{0:T}\in\mathbb{R}^{n(T+1)}$, $\textbf{u}:=u_{0:T-1}\in\mathbb{R}^{n_{i}T}$, $\textbf{w}:=w_{0:T-1}\in\mathbb{R}^{nT}$, and $\textbf{y}:=y_{0:T},\textbf{e}:=e_{0:T-1}\in\mathbb{R}^{n_{o}T}$. We assume there exist known, compact sets $W\subset\mathbb{R}^{n}$ and $E\subset\mathbb{R}^{n_{o}}$ such that $w_{t}\in W$ and $e_{t}\in E$ for each $t\in[T]$. Note that $W$ and $E$ can be estimated from data, e.g., via [14].
@@
To ease exposition, unless noted otherwise, we assume linear time-varying (LTV) dynamics, i.e., $f_{t}(x_{t},u_{t})=A_{t}x_{t}+B_{t}u_{t}$. We add clarifying remarks when a result holds for LTV but not for general nonlinear dynamics (Remarks 1, 2, 6).
@@
Let $C_{k}$, $C_{\urcorner{k}}(\theta^{\star})\subset\mathbb{R}^{n(T+1)+n_{i}T}$ respectively denote a known and an unknown, parameterized constraint set imposed on the state and control trajectory $(\textbf{x},\textbf{u})$, where $\theta^{\star}\in\mathbb{R}^{d}$ denotes a ground truth parameter vector a priori unknown to the constraint learner. Concretely, $C_{k}$ and $C_{\urcorner{k}}(\theta^{\star})$ are respectively defined by inequality constraints $g_{\beta,k}:\mathbb{R}^{n(T+1)+n_{i}T}\rightarrow\mathbb{R}^{n_{k}},\forall\beta\in[n_{k}]$, and $g_{\beta,{\urcorner{k}}}:\mathbb{R}^{n(T+1)+n_{i}T}\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{n_{\urcorner{k}}},\forall\beta\in[n_{\urcorner{k}}]$, as shown below:
$C_{k}$ $:=\{(\textbf{x},\textbf{u}):g_{\beta,k}(\textbf{x},\textbf{u})\leq 0,\forall\beta\in[n_{k}]\},$
@@
$C_{\urcorner{k}}(\theta^{\star})$ $:=\{(\textbf{x},\textbf{u}):g_{\beta,{\urcorner{k}}}(\textbf{x},\textbf{u},\theta^{\star})\leq 0,\forall\beta\in[n_{\urcorner{k}}]\},$
@@
For each $t\in[0,T-1]$, let $\pi_{t}:\mathbb{R}^{n_{o}(t+1)}\rightarrow\mathbb{R}^{n_{i}}$ be a causal output feedback controller which maps realized outputs $y_{0,t}$ to controls $u_{t}$, and let $\pi(\textbf{y}):=(\pi_{0}(y_{0}),\cdots,\pi_{T-1}(y_{0:T-1}))$. We define $\Pi:=\Pi(W,E,C_{k},C_{\urcorner{k}},\theta^{\star})$ to be the set of all causal output feedback laws $\pi(\cdot)$ which generate trajectories that robustly satisfy the constraint sets $C_{k}$ and $C_{\urcorner{k}}(\theta^{\star})$ despite worst-case noise realizations in $W$ and $E$. We then define a cost map $J:\mathbb{R}^{(n+1)T}\times\mathbb{R}^{mT}\rightarrow\mathbb{R}$ and formulate the following robust control problem, which is generally computationally intractable due to the optimization over $\pi(\cdot)$:
@@
$\pi(\cdot)\in\Pi(W,E,C_{k},C_{\urcorner{k}},\theta^{\star}),$ (3d)
II-B System Level Synthesis (SLS) for Output Feedback
A common method for (3) is to find (i) nominal state and control trajectories $(\textbf{z},\textbf{v})\in\mathbb{R}^{n(T+1)+n_{i}T}$ that minimize cost in the noise-free case and (ii) linear output feedback controllers to reject noise [15, 12, 16]. Equivalently, we assume each output feedback law $\pi_{t}(\cdot)$ for $t\in[0,T-1]$ has the form:
@@ II-B System Level Synthesis (SLS) for Output Feedback
where $K_{t,\tau}\in\mathbb{R}^{n_{i}\times n_{o}}$, $\forall t\in[0,T-1]$, $\tau\in[0,t]$. To write (4) compactly, let $\mathcal{A}:=\text{diag}\{A_{0},\cdots,A_{T-1}\}\in\mathbb{R}^{nT\times nT}$, $\mathcal{B}:=\text{diag}\{B_{0},\cdots,B_{T-1}\}\in\mathbb{R}^{nT\times n_{i}T}$, $\mathcal{C}:=\text{diag}\{C_{0},\cdots,C_{T-1}\}\in\mathbb{R}^{n_{o}T\times nT}$, and define $\mathcal{K}\in\mathbb{R}^{n_{i}T\times n_{o}T}$ as follows:
$\mathcal{K}$ $:=\begin{bmatrix}K_{0,0}&O&\cdots&O\\[-7.0pt] K_{1,0}&K_{1,1}&\ddots&O\\[-3.0pt] \vdots&\vdots&\ddots&\vdots\\[-7.0pt] K_{T-1,0}&K_{T-1,1}&\ddots&K_{T-1,T-1}.\end{bmatrix}$ (5)
@@
The policy optimization in (3) can now be recast as an optimization over the nominal state-control trajectory $(\textbf{z},\textbf{v})$ and the output error feedback gain $\mathcal{K}$ in (5). These two optimizations may be decoupled, as in methods based on control contraction metrics (CCM) [16], or coupled, as in system-level synthesis (SLS)-based robust control designs [15, 12].
@@
$\Delta u_{t}$ $=\sum_{\tau=0}^{t}K_{t,\tau}(C_{\tau}\Delta x_{\tau}+e_{\tau}),$ $\forall t\in[0,T-1],$ (7b)
@@
$\mathcal{Z}$ $:=\begin{bmatrix}O_{n\times n(T-1)}&O_{n\times n}\\ I_{n(T-1)}&O_{n(T-1)\times n}\end{bmatrix}\in\mathbb{R}^{nT\times nT}.$ (8)
@@
where the system response $\Phi\in\mathbb{R}^{(n(T+1)+n_{i}T)\times(n+n_{o})T}$ is block-wise defined from $\mathcal{Z},\mathcal{A},\mathcal{B},\mathcal{C}$, and $\mathcal{K}$, as follows:
@@
[15, Prop. 1] If there exists a lower triangular block matrix $\mathcal{K}$ satisfying (9), then $\Phi_{xw}$, $\Phi_{xe}$, $\Phi_{uw}$, and $\Phi_{ue}$, as computed by (11) satisfy the affine equalities:
$\begin{bmatrix}I-Z\mathcal{A}&-Z\mathcal{B}\end{bmatrix}\Phi$ $=\begin{bmatrix}I&O\end{bmatrix}$, (13a)
$\Phi\begin{bmatrix}I-Z\mathcal{A}\\ -\mathcal{C}\end{bmatrix}$ $=\begin{bmatrix}I\\ O\end{bmatrix}$, (13b)
$\Phi_{xw},\Phi_{xe},\Phi_{uw},\Phi_{ue}$ block lower-triangular as in (5) (13c)
Conversely, if $\Phi_{xw}$, $\Phi_{xe}$, $\Phi_{uw}$, and $\Phi_{ue}$ satisfy (13), then $\mathcal{K}$, as computed by (12), is block lower-triangular.
@@
[…]
@@
We then stack constraints $\tilde{g}_{\beta,k}$ (resp., $\tilde{g}_{\beta,{\urcorner{k}}}$) across indices $\beta$ to form the vector-valued maps $\tilde{\textbf{g}}_{k}$ (resp., $\tilde{\textbf{g}}_{\urcorner{k}}$). We also define $\tilde{\textbf{h}}$ to capture the dynamics $z_{t+1}=A_{t}z_{t}+B_{t}v_{t}\ \forall t\in[T]$, as well as (13), as $\tilde{\textbf{h}}(\cdot)=0$. We then pose the following optimization program, which presents a feasible solution to (3) [12, Prop. 4]:
$\min_{\textbf{z},\textbf{v},\Phi}\hskip 14.22636pt$ $\tilde{J}(\textbf{z},\textbf{v},\Phi)$ (14a)
@@
If the dynamics (1a) are nonlinear, then $\tilde{\textbf{h}}$ must be defined to encode $z_{t+1}=f_{t}(z_{t},u_{t})$ for all $t\in[0,T]$, as well as (13), with each $A_{t},B_{t}$ replaced by the Jacobian linearization of $f_{t}$. Moreover, $\tilde{\textbf{g}}_{k}$ and $\tilde{\textbf{g}}_{\urcorner{k}}$ must be defined while accounting for the linearization error related to each $f_{t}$.

### v2

@@ Abstract
We present a method for learning unknown parametric constraints from locally-optimal input-output trajectory data. We assume the data is generated by rollouts of stochastic nonlinear dynamics, under a single state or output feedback law and initial condition but distinct noise realizations, to robustly satisfy underlying constraints despite worst-case noise outcomes. We encode the Karush-Kuhn-Tucker (KKT) conditions of this robust optimal feedback control problem within a feasibility problem to recover constraints consistent with the local optimality of the demonstrations. We prove that our constraint learning method (i) accurately recovers the demonstrator’s policy, and (ii) conservatively estimates the set of policies that ensure constraint satisfaction despite worst-case noise realizations. Moreover, we perform sensitivity analysis, proving that when demonstrations are corrupted by transmission error, the inaccuracy in the learned feedback law scales linearly in the error magnitude. Empirically, our method accurately recovers unknown constraints from simulated noisy, closed-loop demonstrations generated using dynamics, both linear and nonlinear, (e.g., unicycle and quadrotor) and a range of feedback mechanisms.
{IEEEkeywords}
Learning from demonstration, robot safety, robust and optimal control.
1 Introduction
[…]
@@ 1 Introduction
In safety-critical robotics, autonomous agents must infer and enforce hard constraints to guarantee safe operation, such as for collision avoidance in crowded environments. Learning from demonstration (LfD) enables robots to recover unknown parametric constraints from trajectory data [4, 12, 2]. The key idea is that locally-optimal demonstrations generated from a constrained optimal control problem satisfy an associated set of Karush-Kuhn-Tucker (KKT) conditions, which can be inverted to recover constraints via inverse optimal control (IOC). Prior work has shown this is possible for open-loop state and control demonstrations under deterministic dynamics [4]. In practice, however, agents operate with state uncertainty and stochastic dynamics. Thus, instead of open-loop plans, a demonstrator must apply an output feedback control policy to partial, noisy observations. Existing methods cannot recover constraints from such input-output data.
@@
To address this gap, we make the following contributions. We present a novel method to infer a priori unknown parametric constraints from input-output data generated under linear time-varying (LTV) stochastic dynamics with partial observations and output noise. Our method extends the IOC-based constraint learning method in [4] to the partial and noisy observation setting (Sec. 3). We evaluate our method on linear and nonlinear (e.g., unicycle, quadrotor) dynamics and a variety of feedback controllers, successfully recovering constraints from the generated demonstrations (Sec. 4).
@@
Our work draws from and contributes to IOC-based LfD for cost and constraint inference. Specifically, [6, 12] introduce IOC methods assuming known constraints, while [16, 14, 13, 3] (resp., [4, 11]) enable cost (resp., constraint) inference via IOC. Unlike [11], which addresses stochastic demonstrations in discrete spaces, our method recovers constraints with provable accuracy guarantees in continuous space. Also, while [4] use deterministic trajectories, we learn constraints from demonstrations generated via output feedback with noisy, partial observations. On a technical level, we leverage system-level synthesis (SLS) [1, 8, 9] to encode output feedback policies for input-output demonstrations. Whereas prior SLS work focuses on controller design, we apply SLS to infer constraints from noisy, partially observed data.
@@
Notation: $\forall x_{0},\cdots,x_{T}\in\mathbb{R}^{n}$, let $x_{t_{1}:t_{2}}:=(x_{t_{1}},\cdots,x_{t_{2}})\in\mathbb{R}^{n(t_{2}-t_{1}+1)}$ $\forall t_{1},t_{2}\in\{0,\ldots,T\}:=[0,T]$ with $t_{1}\leq t_{2}$. For $M_{1},\cdots,M_{K}\in\mathbb{R}^{m\times n}$, let $M:=\text{diag}\{M_{1},\cdots,M_{K}\}\in\mathbb{R}^{mK\times nK}$ denote the block diagonal matrix with $M_{k}$ as the $k$-th block, $\forall k\in\{1,\ldots,K\}:=[K]$. We denote the $m\times m$ identity matrix, $m\times n$ zero matrix, and $n\times 1$ zero vector by $I_{m}$, $O_{m\times n}$, and $\textbf{0}_{n}$, respectively, $\forall m,n\in\mathbb{N}$, with subscripts omitted for dimensions clear from context.
2 Problem Formulation
@@
We formulate the robust optimal control problem used to generate demonstration data (Sec. 2.1), present system level synthesis (SLS) for output feedback (Sec. 2.2), discuss data generation (Sec. 2.3), and state our problem (Sec. 2.4).
2.1 Robust Optimal Control Problem Formulation
@@
Consider a stochastic dynamical system $f_{t}:\mathbb{R}^{n}\times\mathbb{R}^{n_{i}}\rightarrow\mathbb{R}^{n}$ with additive dynamics and output noise, and initial condition fixed at $x_{0}=\bar{x}$ as shown below: [^1] Uncertainty in $x_{0}$ can be modeled as a non-zero first component of the dynamics noise, replacing $\textbf{0}_{n}$ (see [8]); we do so in our experiments.
@@
Above, $x_{t}\in\mathbb{R}^{n}$, $u_{t}\in\mathbb{R}^{n_{i}}$, $w_{t}\in\mathbb{R}^{n}$, $y_{t}\in\mathbb{R}^{n_{o}}$, and $e_{t}\in\mathbb{R}^{n_{o}}$ respectively denote the state, input, dynamics noise, output, and output noise at time $t\in[0,T]$. Let $\textbf{x}:=x_{0:T}\in\mathbb{R}^{n(T+1)}$, $\textbf{u}:=u_{0:T}\in\mathbb{R}^{n_{i}(T+1)}$, $\textbf{w}:=(\textbf{0}_{n},w_{0:T-1})\in\mathbb{R}^{n(T+1)}$, and $\textbf{y}:=y_{0:T}\in\mathbb{R}^{n_{o}(T+1)},\textbf{e}:=e_{0:T}\in\mathbb{R}^{n_{o}(T+1)}$. We assume there exist known, compact sets $W\subset\mathbb{R}^{n}$ and $E\subset\mathbb{R}^{n_{o}}$ such that $w_{t}\in W\ \forall t\in[T-1]$ and $e_{t}\in E\ \forall t\in[T]$. Note that $W$ and $E$ can be estimated from data, e.g., via [7].
@@
To ease exposition, unless noted otherwise, we assume linear time-varying (LTV) dynamics, i.e., $f_{t}(x_{t},u_{t})=A_{t}x_{t}+B_{t}u_{t}$. We add clarifying remarks when a result holds for LTV but not for general nonlinear dynamics (Remarks 1, 2, 3.4).
@@
Let $S_{k}$, $S_{\urcorner{k}}(\theta^{\star})\subset{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathbb{R}^{(n+n_{i})(T+1)}}$ respectively denote a known and an unknown, parameterized constraint set imposed on the state-control trajectory $(\textbf{x},\textbf{u})$, where $\theta^{\star}\in\mathbb{R}^{d}$ denotes a ground truth parameter vector a priori unknown to the constraint learner. Concretely, $S_{k}$ and $S_{\urcorner{k}}(\theta^{\star})$ are respectively defined by inequality constraints $g_{\beta,k}:{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathbb{R}^{(n+n_{i})(T+1)}}\rightarrow{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathbb{R}},\forall\beta\in[n_{k}]$, and $g_{\beta,{\urcorner{k}}}:{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathbb{R}^{(n+n_{i})(T+1)}}\times\mathbb{R}^{d}\rightarrow{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathbb{R}},\forall\beta\in[n_{\urcorner{k}}]$, as shown below:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}S_{k}}$ $:=\{(\textbf{x},\textbf{u}):g_{\beta,k}(\textbf{x},\textbf{u})\leq 0,\forall\beta\in[n_{k}]\},$
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}S_{\urcorner{k}}}(\theta^{\star})$ $:=\{(\textbf{x},\textbf{u}):g_{\beta,{\urcorner{k}}}(\textbf{x},\textbf{u},\theta^{\star})\leq 0,\forall\beta\in[n_{\urcorner{k}}]\},$
@@
For each $t\in[0,T-1]$, let $\pi_{t}:\mathbb{R}^{n_{o}(t+1)}\rightarrow\mathbb{R}^{n_{i}}$ be a causal output feedback controller which maps realized outputs $y_{0:t}$ to controls $u_{t}$, and let $\pi(\textbf{y}):=(\pi_{0}(y_{0}),\cdots,\pi_{T-1}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}y_{0:T}}))$. We define $\Pi:=\Pi(W,E,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}S_{k}},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}S_{\urcorner{k}}},\theta^{\star})$ to be the set of all causal output feedback laws $\pi(\cdot)$ which generate trajectories that robustly satisfy the constraint sets $S_{k}$ and $S_{\urcorner{k}}(\theta^{\star})$ despite worst-case noise realizations in $W$ and $E$. We then define a cost map $J:{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathbb{R}^{(n+n_{i})(T+1)}}\rightarrow\mathbb{R}$ and formulate the following robust control problem, which is generally computationally intractable due to the optimization over $\pi(\cdot)$:
@@
$\pi(\cdot)\in\Pi(W,E,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}S_{k}},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}S_{\urcorner{k}}},\theta^{\star}){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}.}$ (3d)
2.2 System Level Synthesis (SLS) for Output Feedback
A common approach used rather than solving (3) is to find (i) nominal state-control trajectories $(\textbf{z},\textbf{v})\in{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathbb{R}^{(n+n_{i})(T+1)}}$ that minimize cost in the noise-free case and (ii) linear output feedback controllers to reject noise [17, 8, 10]. Equivalently, we assume each feedback law $\pi_{t}(\cdot)$ for $t\in[0,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}T}]$ has the form:
@@ 2.2 System Level Synthesis (SLS) for Output Feedback
where $K_{t,\tau}\in\mathbb{R}^{n_{i}\times n_{o}}$, $\forall t\in{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}[0,T]}$, $\tau\in[0,t]$. To write (4) compactly, let $\mathcal{A}:=\text{diag}\{A_{0},\cdots,A_{T-1},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}O_{n\times n}}\}\in\mathbb{R}^{n{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}(T+1)}\times n{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}(T+1)}}$, $\mathcal{B}:=\text{diag}\{B_{0},\cdots,B_{T-1},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}O_{n\times n_{i}}}\}\in\mathbb{R}^{n{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}(T+1)}\times n_{i}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}(T+1)}}$, $\mathcal{C}:=\text{diag}\{C_{0},\cdots,C_{T-1},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}C_{T}}\}\in\mathbb{R}^{n_{o}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}(T+1)}\times n{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}(T+1)}}$, and define $\mathcal{K}\in\mathbb{R}^{n_{i}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}(T+1)}\times n_{o}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}(T+1)}}$ by:
$\mathcal{K}$ $:=\begin{bmatrix}K_{0,0}&O&\cdots&O\\[-7.0pt] K_{1,0}&K_{1,1}&\ddots&O\\[-3.0pt] \vdots&\vdots&\ddots&\vdots\\[-7.0pt] {\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}K_{T,0}}&{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}K_{T,1}}&\ddots&{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}K_{T,T}}\end{bmatrix}$ (5)
@@
The policy optimization in (3) can now be recast over the nominal state-control trajectory $(\textbf{z},\textbf{v})$ and the output error feedback gain $\mathcal{K}$ in (5). These two optimizations may be decoupled, as in methods based on control contraction metrics (CCM) [10], or coupled, as in system-level synthesis (SLS)-based robust control designs [17, 8].
@@
$\Delta u_{t}$ $=\sum_{\tau=0}^{t}K_{t,\tau}(C_{\tau}\Delta x_{\tau}+e_{\tau}),$ $\forall t\in[0,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}T}],$ (7b)
@@
$\mathcal{Z}$ $:=\begin{bmatrix}O_{n\times n{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}T}}&O_{n}\\ I_{n{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}T}}&O_{n{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}T}\times n}\end{bmatrix}\in{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathbb{R}^{n(T+1)\times n(T+1)}}.$ (8)
@@
where the system response $\Phi\in\mathbb{R}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}((n+n_{i})(T+1))\times(n+n_{o})(T+1)}}$ is block-wise defined from $\mathcal{Z},\mathcal{A},\mathcal{B},\mathcal{C}$, and $\mathcal{K}$, as follows:
@@
[17, Prop. 1] If there exists a lower block triangular matrix $\mathcal{K}$ satisfying (9), then $\Phi_{xw}$, $\Phi_{xe}$, $\Phi_{uw}$, and $\Phi_{ue}$, as computed by (11) satisfy the affine equalities:
$\begin{bmatrix}I-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathcal{Z}}\mathcal{A}&-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathcal{Z}}\mathcal{B}\end{bmatrix}\Phi$ $=\begin{bmatrix}I&O\end{bmatrix}$, (13a)
$\Phi\begin{bmatrix}I-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathcal{Z}}\mathcal{A}\\ -\mathcal{C}\end{bmatrix}$ $=\begin{bmatrix}I\\ O\end{bmatrix}$, (13b)
$\Phi_{xw},\Phi_{xe},\Phi_{uw},\Phi_{ue}$ lower block triangular as in (5) (13c)
Conversely, if $\Phi_{xw}$, $\Phi_{xe}$, $\Phi_{uw}$, and $\Phi_{ue}$ satisfy (13), then $\mathcal{K}$, as computed by (12), is lower block triangular.
@@
Using [17, Prop. 1], we modify (3) as an optimization problem over nominal trajectories $(\textbf{z},\textbf{v})$ and system responses $\Phi$. Concretely, given $(\textbf{z},\textbf{v})$ and $\Phi$, we characterize the worst-case constraint realizations via the functions $\tilde{g}_{\beta,k}(\cdot)$, $\forall\beta\in[n_{k}]$, and $\tilde{g}_{\beta,{\urcorner{k}}}(\cdot)$, $\forall\beta\in[n_{\urcorner{k}}]$, defined as:
@@
We stack constraints $\tilde{g}_{\beta,k}$ (resp., $\tilde{g}_{\beta,{\urcorner{k}}}$) across indices $\beta$ to form the vector-valued maps $\tilde{\textbf{g}}_{k}$ (resp., $\tilde{\textbf{g}}_{\urcorner{k}}$), and define $\tilde{\textbf{h}}$ to encode (13) and the dynamics $z_{t+1}=A_{t}z_{t}+B_{t}v_{t}\ \forall t\in[T]$, as $\tilde{\textbf{h}}(\cdot)={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\textbf{0}}$. We then pose the optimization program (14), which presents a feasible solution to (3b)-(3d) [8, Prop. 4]:
$\min_{\textbf{z},\textbf{v},\Phi}\hskip 14.22636pt$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}J(\textbf{z},\textbf{v})}$ (14a)
@@
If the dynamics (1a) are nonlinear, $\tilde{\textbf{h}}$ must be defined to encode $z_{t+1}=f_{t}(z_{t},u_{t}),\forall t\in[0,T]$, as well as (13), with each $A_{t},B_{t}$ replaced by the Jacobian linearization of $f_{t}$. Moreover, we must define $\tilde{\textbf{g}}_{k}$ and $\tilde{\textbf{g}}_{\urcorner{k}}$ while accounting for the linearization error related to each $f_{t}$ [9, Sec. III].